# Task Vector Quantization for Memory-Efficient Model Merging

### v1

@@ Abstract
Model merging enables efficient multi-task models by combining task-specific fine-tuned checkpoints. However, storing multiple task-specific checkpoints requires significant memory, limiting scalability and restricting model merging to larger models and diverse tasks. In this paper, we propose quantizing task vectors (i.e., the difference between pre-trained and fine-tuned checkpoints) instead of quantizing fine-tuned checkpoints. We observe that task vectors exhibit a narrow weight range, enabling low-precision quantization ($\leq 4$ bit) within existing task vector merging frameworks. To further mitigate quantization errors within ultra-low bit precision (e.g., $2$ bit), we introduce Residual Task Vector Quantization, which decomposes the task vector into a base vector and offset component. We allocate bits based on quantization sensitivity, ensuring precision while minimizing error within a memory budget. Experiments on image classification and dense prediction show our method maintains or improves model merging performance while using only 8% of the memory required for full-precision checkpoints.
@@ 1 Introduction
Our comprehensive experiments demonstrate that our method maintains or even improves performance with state-of-the-art merging techniques for image classification and challenging dense prediction (see Fig. 2). More importantly, our quantization method modifies only the checkpoints, meaning that it can be integrated into existing model merging frameworks without any modifications. Using just 8% of the original storage, our quantized checkpoints remain effective for both vision tasks. Especially, dense prediction tasks [^1] These include segmentation, depth estimation, and normal estimation. are crucial for vision applications requiring spatially precise reasoning but has been largely overlooked in prior model merging studies. To address this, we first reproduce existing merging approaches in dense prediction settings and evaluate our quantization method, demonstrating its strong performance on these complex tasks.
@@
  • Our method modifies only checkpoints, ensuring seamless integration into existing model merging frameworks. Extensive experiments across multiple domains, including dense prediction, validate its effectiveness.
@@ 2.1 Model Merging
Task arithmetic [23] construct multi-task models through additive task vectors, which are computed by subtracting the pre-trained model’s parameter from those of the fine-tuned model. Unlike direct interpolation, task vectors enable modular composition of specialized models, making them more flexible for multi-task settings. Sparsification-based merging [55, 61], where only the most significant parameters are retained to reduce interference, often complemented by rescaling strategies to enhance stability before integration. However, these approaches require manual tuning of task vector coefficients, which increases computational cost and limits scalability. To mitigate this issue, AdaMerging [58] determines the coefficients adaptively through test-time training. LiNeS [49] simplifies coefficient selection by increasing them linearly with layer depth. To move beyond fixed coefficients for all inputs, router-based approaches [33, 45] dynamically determine instance-wise merging coefficients. Representation Surgery [57] independently learn task-specific parameters at test time to align the merged model’s output with that of each task’s individual model.
As the number of tasks increases, storing multiple fine-tuned checkpoints becomes a major bottleneck. Despite efforts to address storage issues, existing methods like EMR-Merging [20] and TALL Mask [48] rely solely on task-specific binary masks. In contrast, we introduce a more universally applicable strategy that quantizes each task vector, significantly reducing storage overhead while preserving performance. Crucially, our approach integrates seamlessly with existing task vector-based methods, requiring only minimal modifications.
@@ 3.2 Weight Quantization
where $b$ is the number of bits used for quantization, $\Delta$ is the scaling factor, and $z=-\text{Round}(\frac{\theta_{\min}}{\Delta})$ is the zero-point. The scaling factor $\Delta$ ensures that the range $[\theta_{\min},\theta_{\max}]$ is appropriately mapped to the target integer range, while the zero-point $z$ compensates for any asymmetry in the distribution of $\mathbf{\theta}$. The original values can be reconstructed from the quantized representation through the dequantization step:
@@ 4.1 Observation
where $\Delta$ is the scaling factor introduced in Eq. 1. The key insight is that a narrower weight range $[\theta_{\min},\theta_{\max}]$ leads to a smaller quantization error. Therefore, if we can identify a representation with a reduced dynamic range for storing checkpoints, it would result in a lower quantization error. Based on this, we compare the weight ranges of the fine-tuned checkpoint $\theta^{t}_{ft}$, and its corresponding task vector $\tau_{t}$, as illustrated in Fig. 3. The figure shows that the task vector has a weight range that is an order of magnitude smaller than that of the fine-tuned checkpoint. This trend is consistent across various datasets and architectures.
@@ 4.3 Residual Task Vector Quantization (RTVQ)
3:$\theta_{\text{ft\_avg\_ec}}\leftarrow Q(\text{base\_vector},b_{b})+\theta_{\text{pre}}$
@@
5:$\hat{\tau}_{t}\leftarrow Q(\text{offset\_vector},b_{o})+Q(\text{base\_vector},b_{b})$
@@ 5.1 Experimental Setup
Datasets. We evaluate the effectiveness of our proposed quantization method through multi-task model merging experiments on both image classification and dense prediction tasks. Specifically, we adopt the established protocols from Task Arithmetic [23] and AdaMerging [58] to evaluate 8 image classification tasks [54, 25, 6, 19, 37, 28, 43, 8]. Since merging additional tasks increases the storage requirements due to the growing number of model checkpoints, we further expand our experiments to 14 and ultimately 20 classification tasks [26, 17, 38, 39, 47, 10, 11, 53, 4, 9, 40], following [48]. For dense prediction, we follow the FusionBench [44] with the NYUv2 dataset [42], which contains RGB-D images for 13-class semantic segmentation, depth estimation, and normal estimation.
Models. In our image classification experiments, we utilize Vision Transformers of varying scales: ViT-B/32, ViT-B/16, and ViT-L/14, each initialized with pretrained CLIP weights [41] [^2] https://github.com/mlfoundations/open_clip. To ensure a fair comparison across the 8 classification tasks, we adopt the publicly released model checkpoints from Task Arithmetic [23] [^3] https://github.com/mlfoundations/task_vectors. For our extended experiments with 14 and 20 classification tasks, we directly fine-tune CLIP ViT-B/32 to obtain individual task-specific models. For dense prediction tasks, we employ ResNet-50 [18] as the backbone, initializing with ImageNet pre-trained weights. We then fine-tune this model following FusionBench [44] for segmentation, depth estimation, and normal estimation. A comprehensive description of our experimental configurations is provided in Appendix A.
@@ 6 Conclusion
We address the memory overhead of storing full-precision checkpoints in model merging by introducing Task Vector Quantization, which leverages the naturally narrow weight range of task vectors for effective low-precision quantization. To further reduce quantization errors, we propose Residual Task Vector Quantization, which decomposes task vectors into base and residual components, ensuring higher precision where needed. Extensive experiments on image classification and dense prediction tasks show that our method maintains or even improves performance while reducing storage by up to 92%. More importantly, our approach requires no architectural changes, making it seamlessly compatible with existing model merging frameworks. By enabling efficient model merging under limited memory, our work expands the practicality of merging-based multi-task learning across various real-world applications.
Appendix
@@
  • Appendix A: Description of the datasets used in our experiments, the merging method baselines.
@@ A.1 Employed datasets
We evaluated our approach on a diverse set of datasets spanning image classification and dense prediction tasks. For classification, we used SUN397 [54], Cars [25], RESISC45 [6], EuroSAT [19], SVHN [37], GTSRB [43], MNIST [28], DTD [8], CIFAR-10/100 [26], FER2013 [17], Flowers102 [38], Oxford-IIIT Pet [39], PCAM [47], STL10 [10], EMNIST [11], FashionMNIST [53], Food101 [4], KMNIST [9], and Rendered SST-2 [40]. For dense prediction, we employed NYUv2 [42], including segmentation, depth estimation, and normal estimation.
A.2 Merging method baseline
@@ B.1 Impact of quantization
Figure B: Confusion matrices of the cosine similarity among 20 classification task vectors for (a) full-precision (FP32) and (b) 3-bit quantized settings. The diagonal entries are excluded for a clearer comparison.
@@ Appendix C Additional Results
C.1 Full results on 14 and 20 classification tasks
C.2 Additional results for dense prediction tasks
C.3 Comprehensive task-level results
We report the average accuracy across all tasks in Table 1 and Table 2 of the main paper, providing an overall comparison of different quantization methods. In addition, we present the task-specific results for ViT-B/32 and ViT-L/14 in Table E and  F, respectively, to show the performance differences across individual tasks.
C.4 Loss landscape visualizations
We utilized the visualization method introduced in [16]. Using 1,024 test images, we computed the loss values across all 16$\times$16 grid points for each task. For 8 vision tasks, we visualized the loss landscapes of all target task and cross-task pairs using ViT-B/32. The visualization results for cross-tasks can be found in Fig. FGHI, while the visualization results for target tasks are presented in Fig. JK.

### v2

@@ Abstract
Model merging enables efficient multi-task models by combining task-specific fine-tuned checkpoints. However, storing multiple task-specific checkpoints requires significant memory, limiting scalability and restricting model merging to larger models and diverse tasks. In this paper, we propose quantizing task vectors (i.e., the difference between pre-trained and fine-tuned checkpoints) instead of quantizing fine-tuned checkpoints. We observe that task vectors exhibit a narrow weight range, enabling low-precision quantization ($\leq 4$ bit) within existing task vector merging frameworks. To further mitigate quantization errors within ultra-low bit precision (e.g., $2$ bit), we introduce Residual Task Vector Quantization, which decomposes the task vector into a base vector and offset component. We allocate bits based on quantization sensitivity, ensuring precision while minimizing error within a memory budget. Experiments on image classification and dense prediction show our method maintains or improves model merging performance while using only 8% of the memory required for full-precision checkpoints. Our code is available at https://aim-skku.github.io/TVQ/.
@@ 1 Introduction
Our comprehensive experiments demonstrate that our method maintains or even improves performance with state-of-the-art merging techniques for image classification and dense prediction (see Fig. 2). More importantly, our quantization method modifies only the checkpoints, meaning that it can be integrated into existing model merging frameworks without any modifications. Using just 8% of the original storage, our quantized checkpoints remain effective for both vision tasks. Especially, dense prediction tasks [9, 76] [^1] These include segmentation, depth estimation, and normal estimation. are crucial for vision applications requiring spatially precise reasoning but has been largely overlooked in prior model merging studies. To address this, we first reproduce existing merging approaches in dense prediction settings and evaluate our quantization method, demonstrating its strong performance on these complex tasks.
@@
  • Our method modifies only checkpoints, ensuring seamless integration into existing model merging frameworks. Extensive experiments across multiple domains, including dense prediction and natural language processing, validate its effectiveness.
@@ 2.1 Model Merging
Task Arithmetic [28] construct multi-task models through additive task vectors, which are computed by subtracting the pre-trained model’s parameter from those of the fine-tuned model. Unlike direct interpolation, task vectors enable modular composition of specialized models, making them more flexible for multi-task settings. Sparsification-based merging [67, 73], where only the most significant parameters are retained to reduce interference, often complemented by rescaling strategies to enhance stability before integration. However, these methods require manual tuning of task vector coefficients, which increases computational cost and limits scalability. To mitigate this issue, AdaMerging [70] determines the coefficients adaptively through test-time training. LiNeS [59] simplifies coefficient selection by increasing them linearly with layer depth. To move beyond fixed coefficients for all inputs, router-based approaches [40, 54] dynamically compute instance-wise coefficients. Representation Surgery [69] further learns task-specific parameters at test time to align the merged model’s output with each individual task model.
As the number of tasks increases, storing multiple fine-tuned checkpoints becomes a major bottleneck. Despite efforts to address storage issues, existing methods like TALL Mask [58] and TSV-C [19] respectively rely on task-specific binary masks and perform SVD on the task matrix. In contrast, we introduce a more universally applicable strategy that quantizes each task vector, significantly reducing storage overhead while preserving performance. Crucially, our approach integrates seamlessly with existing task vector-based methods, requiring only minimal modifications.
@@ 3.2 Weight Quantization
where $b$ is the number of bits used for quantization, $\Delta$ is the scaling factor, and $z=-\text{Round}(\frac{\theta_{\min}}{\Delta})$ is the zero-point. The scaling factor $\Delta$ ensures that the range $\theta_{\min},\theta_{\max}]$ is appropriately mapped to the target integer range, while the zero-point $z$ compensates for any asymmetry in the distribution of $\mathbf{\theta}$. The original values can be reconstructed from the quantized representation through the dequantization step:
@@ 4.1 Observation
where $\Delta$ is the scaling factor introduced in Eq. 1. The key insight is that a narrower weight range $\theta_{\min},\theta_{\max}]$ leads to a smaller quantization error. Therefore, if we can identify a representation with a reduced dynamic range for storing checkpoints, it would result in a lower quantization error. Based on this, we compare the weight ranges of the fine-tuned checkpoint $\theta^{t}_{ft}$, and its corresponding task vector $\tau_{t}$, as illustrated in Fig. 3. The figure shows that the task vector has a weight range that is an order of magnitude smaller than that of the fine-tuned checkpoint. This trend is consistent across various datasets and architectures.
@@ 4.3 Residual Task Vector Quantization (RTVQ)
[…]
@@ 5.1 Experimental Setup
Datasets. We evaluate our method through a series of multi-task model merging experiments. For image classification, we adopt the 8-task protocol from [28] and expand to 14 and 20 tasks [58] to challenge the scalability of our method against increasing storage requirements. For dense prediction, we follow [53] with the NYUv2 [50]. For NLP tasks, we use the GLUE benchmark [57] with the same setup in [25]. Additional details are in the Appendix A.
Models. In our image classification experiments, we utilize CLIP [48]’s visual encoder ViT-B/32 and ViT-L/14. For dense prediction tasks, we employ ResNet-50 [23] as the backbone and follow the protocol of [53] for each task. For NLP tasks, we use RoBERTa-base [36], following the experimental setup of [25]. See Appendix A for details.
@@ 5.3 Analyses
Comparison with prior compression methods. To demonstrate the effectiveness of our method, we compare against prior approaches such as TALL-Mask [58] and TSV-C [19]. Following the TALL-Mask protocol, we evaluate both storage and accuracy. As shown in Table 6, our 3-bit TVQ achieves the highest compression among all methods. It also supports flexible compression via adjustable bitwidths. When applied to the singular components of TSV-C, our quantization yields additional savings with minimal accuracy loss.
[…]
TALL Mask 95.7 23.1
TSV-C 95.7 20.7
Ours 95.9 18.3
TSV-C + Ours 95.7 11.9
Table 6: Comparison on ViT-B/32 across 8 tasks.
Figure 11: NLP tasks analysis.
Task generalizability of the proposed method. To assess the generality of our method, we conducted additional experiments on NLP tasks (Fig. 11). Remarkably, our approach maintains stable and competitive performance, closely matching the full-precision baseline even at 3-bit quantization. This demonstrates the robustness and adaptability of our method across diverse tasks and compression levels.
@@ 6 Conclusion
We address the memory overhead in model merging by introducing Task Vector Quantization, which exploits the narrow weight range of task vectors for effective low-precision storage. To further reduce errors, we propose Residual Task Vector Quantization, decomposing vectors into base and residual parts for higher accuracy where needed. Experiments on various tasks show that our method preserves or improves performance while reducing storage by up to 92% without any architectural changes.
Acknowledgment
This research was partly supported by the MSIT(Ministry of Science and ICT), Korea, under the Graduate School of Metaverse Convergence support program(IITP-2025-RS-2023-00254129) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation). This research was partly supported by the MSIT(Ministry of Science, ICT), Korea, under the Global Research Support Program in the Digital Field program(RS-2024-00425354) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation). This research was partly supported by the MSIT(Ministry of Science and ICT), Korea, under the ICAN(ICT Challenge and Advanced Network of HRD) support program(RS-2024-00436934) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation).
Supplementary Material:
Task Vector Quantization for Memory-Efficient Model Merging
Youngeun Kim1*  Seunghwan Lee2*  Aecheon Jung2*  Bogon Ryu2  Sungeun Hong2†
1Yale University  2Sungkyunkwan University
youngeun.kim@yale.edu   {simon2, kasurashan, bogon.ryu, csehong}@skku.edu
@@ Acknowledgment
  • Appendix A: Details of the datasets, model checkpoints, and merging methods used in our experiments.
@@ A.1 Employed datasets
We evaluated our approach on a broad and diverse set of datasets that span three major categories: image classification, dense prediction, and natural language processing.
For image classification, we conducted experiments on a total of 19 datasets that cover a wide range of domains and visual characteristics. Specifically, we used SUN397 [66] for scene recognition, Cars [31] for fine-grained vehicle classification, and RESISC45 [7] and EuroSAT [24] for remote sensing imagery. We also included digit and character recognition datasets such as SVHN [44], MNIST [34], EMNIST [13], FashionMNIST [65], and KMNIST [11]. Additional benchmarks included GTSRB [52] for traffic sign recognition, DTD [10] for texture classification, CIFAR-10/100 [32] and STL10 [12] for general object recognition, FER2013 [22] for facial expression recognition, Flowers102 [45] and Oxford-IIIT Pet [46] for fine-grained species classification, PCAM [56] for histopathology image analysis, Food101 [4] for food recognition, and Rendered SST-2 [47], which contains rendered images generated from sentiment classification data.
For dense prediction tasks, we used NYUv2 [50], which provides RGB-D indoor images annotated for multiple tasks, including 13-class semantic segmentation, depth estimation, and surface normal estimation. This dataset was chosen to evaluate the ability of our method to handle multi-modal dense prediction problems.
Finally, for NLP tasks, we adopted the GLUE benchmark [57], which consists of multiple language understanding tasks. We reported the Matthews correlation coefficient for CoLA [60], Pearson and Spearman correlations for STS-B [6], and accuracy for the remaining tasks, including SST-2 [51], MRPC [15], QQP [29], MNLI [62], QNLI [49], and RTE [21]. These benchmarks comprehensively evaluate our method across different input modalities and task types.
A.2 Model checkpoints
For our experiments, we address three distinct domains. In image classification, we utilize Vision Transformers of varying scales: ViT-B/32 and ViT-L/14, initialized with pretrained CLIP weights [48] [^3] https://github.com/mlfoundations/open_clip. To ensure a fair comparison across the 8 classification tasks, we adopt the publicly released model checkpoints from Task Arithmetic [28] [^4] https://github.com/mlfoundations/task_vectors. For our extended experiments with 14 and 20 classification tasks, we fine-tune CLIP ViT-B/32 following Tall-Mask [58] to obtain task-specific models. For dense prediction tasks, we employ ResNet-50 [23] as the backbone, initializing with ImageNet pre-trained weights. We then fine-tune this model following FusionBench [53] for segmentation, depth estimation, and normal estimation. For NLP tasks, we use RoBERTa-base [36] with the publicly available fine-tuned weights from EMR-Merging [25] [^5] https://github.com/harveyhuang18/EMR_Merging.
A.3 Merging method baseline
@@
TSV-Compress [19] exploits the low-rank structure of layer-wise task matrices using singular value decomposition, enabling effective compression of task vectors.
@@ B.1 Impact of quantization
Figure B: Confusion matrices of the cosine similarity among 20 classification task vectors for (a) full-precision (FP32) and (b) 3-bit quantized settings. The diagonal entries are excluded for a clearer comparison. Indices 0 to 19, respectively, correspond to: MNIST, Cars, DTD, EuroSAT, GTSRB, RESISC45, SUN397, SVHN, PCAM, CIFAR100, STL10, OxfordIIITPet, Flowers102, FER2013, CIFAR10, Food101, RenderedSST2, EMNIST, FashionMNIST, and KMNIST.
@@ Appendix C Additional Results
C.1 Full results on 14 and 20 classification tasks
C.2 Additional results for dense prediction tasks
C.3 Comprehensive task-level results
We report the average accuracy across all tasks in Table 1 and Table 2 of the main paper. This provides an overall comparison of different quantization methods and shows how well each method generalizes across multiple tasks. We also present detailed task-specific results for ViT-B/32 and ViT-L/14 in Table E and  F. These results enable closer inspection of per-task behavior and reveal where certain methods perform well or poorly. Such analysis is important for understanding how sensitive quantization strategies are to different task characteristics.
C.4 Loss landscape visualizations
We utilized the visualization method introduced in [20]. Using 1,024 test images, we computed the loss values across all $\times16$ grid points for each task. For 8 vision tasks, we visualized the loss landscapes of all target task and cross-task pairs using ViT-B/32. The visualization results for cross-tasks can be found in Fig. FGHI, while the visualization results for target tasks are presented in Fig. JK.